which of the expressions below are equal to


"3\u20224\u20228 - 5\u20226\u202216 + 7\u20228\u202232 - ..."


when summed to the "k" terms?


1)

"k""+ 1"

"\u03a3" "a(a+1)(-1)" a+1 "2^a"

"a =" "3"



2)

"k"

"\u03a3" "(2a + 1) (2a + 2) 2" a + 2

"a =" "1"



3)

"k""+ 1"

"\u03a3" "4(-1)^a a(2a-1)2^a"

"a =" "2"



4)


"k"

"\u03a3" "(-1)^a a(a+1)(a+5)"

"a =" "3"

 




5) none of the above


1
Expert's answer
2021-06-07T19:27:39-0400

Given, the expression

3•4•8−5•6•16+7•8•32−... (upto k times).

1)

"Σ_{a=3}^{k+1} a(a+1)(-1)^{a+1}2^a\\
=3.4.8-4.5.16+5.6.32-...(\text{upto k +1 terms})"

2)

"Σ_{a=1}^{k} (2a+1)(2a+2)2^{a+2}\\
=3.4.8+5.6.16+7.8.32+...(\text{upto k terms})"

3)

"Σ_{a=2}^{k+1}4(-1)^a a(2a-1)2^{a}\\
=4.2.3.4-4.3.5.8+4.4.7.16+...(\text{upto k +1 terms})"

4)

"Σ_{a=3}^{k}(-1)^a a(a+1)(a+5)\\
=-3.4.8+4.5.9-5.6.10+...(\text{upto k terms})\\
\text{Since, none of the above series matched woth the given series.}\\\text{
Therefore, correct option is (5) none of these.}"